# SceneSplat: Gaussian Splatting-based Scene Understanding with Vision-Language Pretraining

### v1

@@ A.4 Self-Distillation Representation Learning
In self-distilled representation learning, we adopt the framework of [52], combining DINO loss and coding rate regularization on pooled encoder features. We configure a batch size of 24 and set the grid sampling resolution to 0.02 for partitioning Gaussian scenes. After grid sampling, scenes are randomly partitioned into base views $\mathcal{G}_{b}$, which are then used to generate global and local crops. For global crops, we randomly select a Gaussian splat from $\mathcal{G}_{b}$ and sample its $K$ neighbors, where $K\sim[0.4,1.0]\times 256,000$. For local crops, we sample $K\sim[0.1,0.4]\times 102,400$. Each batch includes $N_{g}=2$ global views and $N_{l}=3$ local views to balance scene coverage and granularity. For the DINO loss $\mathcal{L}_{\text{DINO}}$, we extract tokenized features from the student encoder $f_{\theta}(\cdot)$ and teacher encoder $f_{\phi}(\cdot)$, where $\phi$ is updated via exponential moving average (EMA). A global representation $\bar{z}$ is obtained by mean pooling over spatial tokens. This representation is projected into a latent space using a 3-layer MLP $P_{\text{DINO}}$ (dimensions: 2048 → 2048 → 256) followed by $\ell_{2}$-normalization. Then we calculate similarity loss:
@@
For regularization, we use the negative of coding rate:
@@
$R_{\epsilon}$ approximates the rate-distortion function of a Gaussian random variable with covariance $\Gamma$, becoming exact as $\epsilon\to 0$. More specifically, it quantifies the spread of the covariance, even when the underlying variables are not strictly Gaussian.
For the iBOT loss, we mask 50% of global views with masking ratios sampled uniformly from $r\sim[0.2,0.7]$. Masked tokens are replaced by a learnable token ($T_{\text{mask}}$). Note here the masked global view embedding feature goes through the student network and outputs $\hat{T^{\prime}}(\theta_{s})\in\mathbb{R}^{N^{\prime}\times d_{\text{r}}}$ while $d_{\text{r}}$ note the representation dimension. and the unmasked global view embedding forward to teacher network and output $\hat{T^{\prime}}(\theta_{t})\in\mathbb{R}^{N^{\prime}\times d_{\text{r}}}$. We detach the teacher’s output and calculate the simple iBOT loss using 3-layer MLP $P_{\text{iBOT}}(\cdot)$ (dimensions: 256 → 256 → 32) as projector:
@@
Only the masked regions $M$ contribute to the iBOT loss calculation, Following [52], we simplify the original iBOT implementation by removing the centering operation and online tokenizer, replacing them with a pairwise cosine-similarity loss for feature alignment. When MGM loss is enabled then for one global view we will employ less aggressive augmentation and this view will contribute to both iBOT loss and MGM loss.
@@ A.5 Autoencoder for feature compression
Building on the idea from [36], we train a scene-specific language autoencoder on precomputed vision-language embeddings. This model compresses high-dimensional SigLIP features into a low-dimensional latent space, enabling efficient storage of language features within Gaussian representations. The encoder (by default 5-layer architecture: [384, 192, 96, 48, 16]) generates compact latent codes, while the symmetric decoder (by default 5-layer architecture: [48, 96, 192, 384, 768]) reconstructs the original high-dimensional embeddings. This design reduces memory overhead while preserving semantic fidelity. Unlike [36], our model is trained on 3D Gaussian features rather than 2D image-level data. For detailed ablations on the autoencoder architecture and training paradigm, see Tab. B.
@@
We follow the same loss functions in Eqs. 2 and 3 for feature alignment. When combining MGM with language alignment, we adhere to the MGM process by masking input tokens and applying the compressed language alignment loss only to the masked regions. Given the comparable magnitudes of all loss terms, we assign equal weights ($\omega_{\text{MGM}}=\omega_{\text{DINO}}=\omega_{\text{iBOT}}=\omega_{\text{LA}}=1.0$) to maintain the balance of different objective during training.
B Further Experimental Results
B.1 SceneSplat Zero-shot Segmentation
Figure A: Comparison of SceneSplat Zero-Shot Versus Fully Supervised Segmentation Across Object Classes. Notably, our zero-shot segmentation after vision language pretraining demonstrates better performance for 18 object classes, predominantly small objects such as headphones, printer, pot, and clothes hanger.
Zero-shot vs. Fully Supervised. Fig. A compares the class-wise zero-shot 3D segmentation results with those of the current state-of-the-art supervised method [51]. Points above the diagonal line (colored red) represent classes where zero-shot segmentation outperforms fully supervised approaches, while points below (colored blue) indicate the opposite. Our zero-shot segmentation achieves superior performance for 18 object classes in the ScanNet++ benchmark, predominantly small objects such as headphones, printers, pots, and clothes hangers. These results demonstrate the robust prior knowledge acquired by our model through vision-language pretraining.
Consistency Issue in Label Collection. Tab. A presents our zero-shot segmentation results on the ScanNet20 benchmark. We observe that performance on this particular benchmark is significantly lower compared to the state-of-the-art results we achieve on the other three benchmarks. Through detailed analysis, we identify that the issue stems from inconsistencies in the 3DGS language label collection process. Specifically, the SAMv2+SigLIP2 pipeline we employ does not guarantee temporal consistency, especially for large background objects such as walls and floors, as illustrated in Fig. B. This inconsistency results in corrupted feature fields for Gaussians associated with these regions, subsequently leading to reduced zero-shot segmentation performance. Addressing this temporal consistency issue remains an important direction for our future work.
Method Training Source ScanNet20 (20)
f-mIoU f-mAcc
OpenScene ScanNet 57.5 72.4
Mosaic3D ScanNet 65.0 82.5
PLA ScanNet 19.1 41.5
RegionPLC ScanNet 55.6 76.3
OV3D ScanNet 64.0 76.3
SceneSplat ScanNet 35.4 57.9
Table A: Zero-Shot 3D Semantic Segmentation on ScanNet20 Benchmark. The results for the baselines are taken from [21]. Ours observes many faulty predictions and have poor performance, we identify the issue in the inconsistency during 2D feature map collection when labeling Gaussians, which leads to corrupted 3DGS-feature pairs.
Refer to caption Refer to caption
Figure B: Consistency Issue During 2D Vision-Language Feature Map Collection on ScanNet. Erroneous 2D feature maps are collected shown by the mislabeled regions in the neighboring figures. The root cause is that the SAMv2+SigLip2 process we use does not guarantee temporal consistency.
B.2 Gaussian Self-Supervised Pretraining
Method Scannet++ 2D Scannet++ 3D
L2 cosine L2 cosine
VAE16 layer5 0.008 0.182 0.005 0.028
VAE16 layer6 0.008 0.172 0.005 0.023
VAE64 layer5 0.007 0.081 0.004 0.014
VAE64 layer6 0.007 0.079 0.004 0.013
Table B: AutoEncoder Ablation Experiments. We report the feature compression performance with $\mathcal{L}_{2}$ loss and cosine similarity on unseen 3D language label. We ablate on different autoencoder architectures and training sources.
In Tab. B, we present an ablation study comparing various autoencoder architectures trained on the ScanNet++ training set and evaluated on the validation set. In the “Scannet++ 2D” columns, the autoencoder is trained using SigLIP2 features preprocessed from images, while in the “Scannet++ 3D” columns, it is trained directly on Gaussian SigLIP2 features. At inference, we measure both the L2 distance and cosine similarity on 3D Gaussian SigLIP2 features. The results indicate that training the autoencoder directly on 3D data yields notably better performance, whereas increasing the network depth from layer5 to layer6 leads to only marginal improvements. Moreover, a 64-dimensional latent space consistently outperforms the 16-dimensional counterpart. Consequently, and as detailed in Sec. A.5, we adopt the 3D-trained autoencoder with layer5 and a 16-dimensional latent space as our default configuration, striking a favorable balance between efficiency and accuracy.
Method ScanNet20 (20)
mIoU mAcc
LA16 76.3 83.9
LA64 75.8 82.2
LA16 (MGM) 76.2 83.7
Table C: Language Alignment Loss Ablation. We conduct an ablation study on low-dimensional language feature compression and Masked Gaussian Modeling (MGM)-based pretraining to evaluate their impact on semantic segmentation performance.
We evaluate the impact of language feature alignment loss dimensionality (64 vs.16) when pretraining exclusively on ScanNet++ and testing on the ScanNet20 benchmark. As shown in Tab. C, We implement two 3-layer MLP architectures: one with uniform dimensions (64 → 64 → 64) for 64 dimension language feature and another with progressively reduced dimensions (64 → 32 → 16). We observe that reducing the latent dimension from 64D to 16D improves mIoU by +$0.5\%$ in mIoU and +0.7% in mAcc. This suggests that lower-dimensional language features with hierarchical dimensionality reduction preserve critical indoor scene semantics information.
We visualize pretraining results using Masked Gaussian Modeling (MGM) and language feature alignment loss in
. The model effectively reconstructs masked Gaussian parameters (e.g., position, scale, rotation) and predicts semantically meaningful language-aligned features for indoor scene understanding.
Tab. D evaluates how masking ratio and grid size in Masked Gaussian Modeling (MGM) affect downstream semantic segmentation performance. Key observations include: (1) Small Masking ratios (0.4) degrade mIoU by $0.9\%$ due to insufficient context learning compared with the 0.6. Using a larger grid size (0.15m) for masking reduces fine-grained detail recovery, leading to a 0.5% drop in mIoU compared to a 0.05m grid size.
Mask Ratio ScanNet20 Mask ScanNet20
mIoU mAcc Size mIoU mAcc
0.4 76.1 83.6 0.02 76.9 84.5
0.5 76.4 84.1 0.05 77.0 84.5
0.6 77.0 84.4 0.10 76.6 83.8
0.7 76.7 83.8 0.15 76.5 84.1
Table D: Masked Gaussian Modeling Ablation Experiment. We analyze the impact of masking ratios and grid sizes in Masked Gaussian Modeling (MGM) on semantic segmentation performance using the ScanNet20 benchmark.
Method ModelNet10 (10) Omniobject3D (83)
Linear MLP-3 Linear MLP-3
MGM 77.3 83.1 50.3 57.5
+DINO 74.5 80.2 40.5 42.4
+iBOT 65.2 73.8 38.4 40.8
+LA 68.1 74.8 26.5 31.2
MGM+LA 83.1 84.6 47.3 53.2
Table E: Cross Domain Linear Probe Experiments. We report the mAcc results in the ablation of GaussianSSL on object-level classification tasks using the linear probe.
To evaluate global scene understanding, we design a cross-domain classification task (Tab. E) using object-level Gaussian splats from [28, 29]. We freeze the encoder and embedding layers, training only (1) A linear probe (final layer) to map features to logits. (2) A 3-layer MLP (512 → 256 → classes) for nonlinear evaluation. We find that adding progressively the DINO, iBOT, and language alignment (LA) losses degrade classification accuracy, By using Masked pretraining and language alignment loss achieves the best performance on ModelNet10 (furniture), while MGM alone excels on OmniObject3D (common objects). In Furniture datasets (e.g., ModelNet10) achieve the best 80% accuracy, while challenging object datasets (e.g., OmniObject3D) peak at  50%. The misalignment with scene-level benchmarks stems from domain gaps: the model maps architecturally distinct objects (e.g., fridge, oven) to similar scene-level semantics (e.g., ”kitchen”).
C Datasets Curation and Statistics
C.1 Analysis
Figure F: Example Incomplete Scenes in ARKitScenes.
Figure G: Blurry Scenes from 3RScan.
C.2 Visualization
Hypersim GS. Ground truth (left) and 3DGS rendering results (right).
D Architecture Designs
@@
Table F: Model Configs for Vision-Language Pretraining.
@@
Table G: Model Configs for GaussianSSL Pretraining and Downstream Semantic Segmentation.
@@
random crop ratio: (0.1, 0.4) max: 256000
@@
Table H: Data Augmentations. Following [31], we design the data augmentation for 3D scenes for global and local views.
@@
We detail our network design choices in Tabs. F and G. For the 3D backbone, we adopt the state-of-the-art Point Transformer [51], enhanced with Flash Attention to substantially improve computational efficiency. To optimize feature embedding for large scale Gaussians, we use sparse convolutions, which preserve geometric details while minimizing memory overhead.
E Dataset License
@@
SceneSplat-7K is constructed from several established indoor datasets, each attached with a specific license: ARKitScenes [1] (Apple Open Source License), Replica [45] (Replica Research License), ScanNet [5] (ScanNet Terms of Use), ScanNet++ [57] (ScanNet++ Terms of Use), Hypersim [42] (CC BY-SA 3.0), 3RScan [48] (CC BY 4.0), and Matterport3D [2] (non-commercial license). For sources that permit redistribution, we plan to release our data on Hugging Face under their respective terms of use. For datasets that require special permission, we will co-host the data only after receiving approval from the original teams.
F Limitations

### v2

@@ A.4 Self-Distillation Representation Learning
In self-distilled representation learning, we adopt the framework of [52], combining DINO loss and coding rate regularization on pooled encoder features. We configure a batch size of 24 and set the grid sampling resolution to 0.02 for partitioning Gaussian scenes. After grid sampling, the scenes are randomly partitioned into base views $\mathcal{G}_{b}$, which are then used to generate global and local crops. For global crops, we randomly select a Gaussian splat from $\mathcal{G}_{b}$ and sample its $K$ neighbors, where $K\sim[0.4,1.0]\times 256,000$. For local crops, we sample $K\sim[0.1,0.4]\times 102,400$. Each batch includes $N_{g}=2$ global views and $N_{l}=3$ local views to balance the scene coverage and granularity. For the DINO loss $\mathcal{L}_{\text{DINO}}$, we extract tokenized features from the student encoder $f_{\theta}(\cdot)$ and teacher encoder $f_{\phi}(\cdot)$, where $\phi$ is updated via exponential moving average (EMA). A global representation $\bar{z}$ is obtained by mean pooling over spatial tokens. This representation is projected into a latent space using a 3-layer MLP $P_{\text{DINO}}$ (dimensions: 2048 → 2048 → 256) followed by $\ell_{2}$-normalization. Then, we calculate the similarity loss:
@@
For regularization, we use the negative of the coding rate:
@@
$R_{\epsilon}$ approximates the rate-distortion function of a Gaussian random variable with covariance $\Gamma$, becoming exact as $\epsilon\to 0$. More specifically, it quantifies the spread of covariance, even when the underlying variables are not strictly Gaussian.
For the iBOT loss, we mask 50% of the global views with masking ratios sampled uniformly from $r\sim[0.2,0.7]$. The masked tokens are replaced by a learnable token ($T_{\text{mask}}$). Note here the masked global view embedding feature goes through the student network and outputs $\hat{T^{\prime}}(\theta_{s})\in\mathbb{R}^{N^{\prime}\times d_{\text{r}}}$, where $d_{\text{r}}$ denotes the representation dimension, while the unmasked global view embedding is forwarded to the teacher network and outputs $\hat{T^{\prime}}(\theta_{t})\in\mathbb{R}^{N^{\prime}\times d_{\text{r}}}$. We detach the teacher’s output and calculate the simple iBOT loss using 3-layer MLP $P_{\text{iBOT}}(\cdot)$ (dimensions: 256 → 256 → 32) as the projector:
@@
Only the masked regions $M$ contribute to the iBOT loss calculation. Following [52], we simplify the original iBOT implementation by removing the centering operation and online tokenizer and replacing them with a pairwise cosine-similarity loss for feature alignment. When MGM loss is enabled, for one global view, we will employ less aggressive augmentation, and this view will contribute to both iBOT loss and MGM loss.
@@ A.5 Autoencoder for Feature Compression
Building on the idea from [36], we train a scene-specific language autoencoder on precomputed vision-language embeddings. This model compresses high-dimensional SigLIP features into a low-dimensional latent space, enabling the efficient storage of language features within Gaussian representations. The encoder (by default 5-layer architecture: [384, 192, 96, 48, 16]) generates compact latent codes, while the symmetric decoder (by default 5-layer architecture: [48, 96, 192, 384, 768]) reconstructs the original high-dimensional embeddings. This design reduces memory overhead while preserving semantic fidelity. Unlike [36], our model is trained on 3D Gaussian features rather than 2D image-level data. For detailed ablations on the autoencoder architecture and training paradigm, see Tab. E.
@@
We follow the same loss functions in Eqs. 2 and 3 for feature alignment. When combining MGM with language alignment, we adhere to the MGM process by masking the input tokens and applying compressed language alignment loss only to the masked regions. Given the comparable magnitudes of all loss terms, we assign equal weights ($\omega_{\text{MGM}}=\omega_{\text{DINO}}=\omega_{\text{iBOT}}=\omega_{\text{LA}}=1.0$) to maintain the balance of different objectives during training.
A.7 Model Architecture
@@
Table A: Model Configs for Vision-Language Pretraining.
@@
Table B: Model Configs for GaussianSSL Pretraining and Downstream Semantic Segmentation.
@@
random crop ratio: (0.1, 0.4) max: 102400
@@
Table C: Data Augmentations. Following [31], we design the data augmentations for 3D scenes for global and local views.
@@
The network design choices are described in detail in Tabs. A and B. For the 3D backbone, we adopt the state-of-the-art Point Transformer [51], enhanced with Flash Attention, to substantially improve computational efficiency. To optimize feature embedding for scene-level Gaussians, we use sparse convolutions, which preserve geometric details while minimizing the memory overhead.
B Further Experimental Results
B.1 SceneSplat Zero-shot Segmentation
Figure A: Comparison of SceneSplat Zero-Shot Versus Fully Supervised Segmentation Across Object Classes. Notably, our zero-shot segmentation after vision language pretraining demonstrates better performance for 18 object classes, predominantly small objects such as headphone, printer, pot, and clothes hanger.
Zero-shot vs. Fully Supervised. Fig. A compares the class-wise zero-shot 3D segmentation results with those of the current state-of-the-art supervised method [51]. Points above the diagonal line (colored red) represent classes where zero-shot segmentation outperforms fully supervised approaches, whereas points below (colored blue) indicate the opposite. Our zero-shot segmentation achieves superior performance for 18 object classes in the ScanNet++ benchmark, predominantly small objects such as headphones, printers, pots, and clothes hangers. These results demonstrate the robust prior knowledge acquired by our model through vision-language pretraining.
Consistency Issue in Label Collection. Tab. D presents our zero-shot segmentation results on the ScanNet20 benchmark. We observe that the performance on this particular benchmark is significantly lower compared to the state-of-the-art results we achieve on the other three benchmarks. Through a detailed analysis, we identify that the issue stems from inconsistencies in the 3DGS language label collection process. Specifically, the SAMv2+SigLIP2 pipeline that we employ does not guarantee temporal consistency, especially for large background objects such as walls and floors, as illustrated in Fig. B. This inconsistency results in corrupted feature fields for Gaussians associated with these regions, subsequently leading to reduced zero-shot segmentation performance. Addressing this temporal consistency issue remains an important direction for future research.
Method Training Source ScanNet20 (20)
f-mIoU f-mAcc
OpenScene ScanNet 57.5 72.4
Mosaic3D ScanNet 65.0 82.5
PLA ScanNet 19.1 41.5
RegionPLC ScanNet 55.6 76.3
OV3D ScanNet 64.0 76.3
SceneSplat ScanNet 35.4 57.9
Table D: Zero-Shot 3D Semantic Segmentation on ScanNet20 Benchmark. The results for the baselines are taken from [21]. Ours observes many faulty predictions and have poor performance, we identify the issue in the inconsistency during 2D feature map collection when labeling Gaussians, which leads to corrupted 3DGS-feature pairs.
Refer to caption Refer to caption
Figure B: Consistency Issue During 2D Vision-Language Feature Map Collection on ScanNet. Erroneous 2D feature maps are collected, as shown by the mislabeled regions in the neighboring figures. The root cause is that the SAMv2+SigLip2 process we use does not guarantee temporal consistency.
B.2 3DGS Self-Supervised Pretraining
Method ScanNet++ 2D ScanNet++ 3D
L2 cosine L2 cosine
VAE16 layer5 0.008 0.182 0.005 0.028
VAE16 layer6 0.008 0.172 0.005 0.023
VAE64 layer5 0.007 0.081 0.004 0.014
VAE64 layer6 0.007 0.079 0.004 0.013
Table E: AutoEncoder Ablation Experiments. We report the feature compression performance with $\mathcal{L}_{2}$ loss and cosine similarity on unseen 3D language label. We ablate on different autoencoder architectures and training sources.
In Tab. E, we present an ablation study that compares various autoencoder architectures trained on the ScanNet++ training set and evaluated on the validation set. In the “ScanNet++ 2D” columns, the autoencoder is trained using SigLIP2 features preprocessed from images, while in the “ScanNet++ 3D” columns, it is trained directly on Gaussian SigLIP2 features. At inference, we measure both the L2 distance and cosine similarity on 3D Gaussian SigLIP2 features. The results indicate that training the autoencoder directly on 3D data yields notably better performance, whereas increasing the network depth from layer5 to layer6 leads to only marginal improvements. Moreover, the 64-dimensional latent space consistently outperforms the 16-dimensional counterpart. Consequently, as detailed in Sec. A.5, we adopt the 3D-trained autoencoder with layer5 and a 16-dimensional latent space as our default configuration, striking a favorable balance between efficiency and accuracy.
Method ScanNet20 (20)
mIoU mAcc
LA16 76.3 83.9
LA64 75.8 82.2
LA16 (MGM) 76.2 83.7
Table F: Language Alignment Loss Ablation. We conduct an ablation study on low-dimensional language feature compression and Masked Gaussian Modeling (MGM)-based pretraining to evaluate their impact on semantic segmentation performance.
We evaluate the impact of language feature alignment loss dimensionality (64 vs.16) when pretraining exclusively on ScanNet++ and testing on the ScanNet20 benchmark. As shown in Tab. F, we implement two 3-layer MLP architectures: one with uniform dimensions (64 → 64 → 64) for 64 dimension language feature and another with progressively reduced dimensions (64 → 32 → 16). We observe that reducing the latent dimension from 64D to 16D improves mIoU by +$0.5\%$ in mIoU and +0.7% in mAcc. This suggests that lower-dimensional language features with hierarchical dimensionality reduction preserve critical indoor scene semantic information.
We visualize the pretraining results using Masked Gaussian Modeling (MGM) and language feature alignment loss in Fig. D. The model effectively reconstructs masked Gaussian parameters (e.g., position, scale, and rotation) and predicts semantically meaningful language-aligned features for indoor scene understanding.
Tab. G evaluates how masking ratio and grid size in Masked Gaussian Modeling (MGM) affect downstream semantic segmentation performance. Key observations include: (1) small masking ratios (0.4) degrade mIoU by $0.9\%$ due to insufficient context learning compared to the ratio 0.6; (2) using a larger grid size $0.15m$ for masking reduces fine-grained detail recovery, leading to a $0.5\%$ drop in mIoU compared to a $0.05m$ grid size.
Mask Ratio ScanNet20 Mask ScanNet20
mIoU mAcc Size mIoU mAcc
0.4 76.1 83.6 0.02 76.9 84.5
0.5 76.4 84.1 0.05 77.0 84.5
0.6 77.0 84.4 0.10 76.6 83.8
0.7 76.7 83.8 0.15 76.5 84.1
Table G: Masked Gaussian Modeling Ablation Experiment. We analyze the impact of masking ratios and grid sizes in Masked Gaussian Modeling (MGM) on semantic segmentation performance using the ScanNet20 benchmark.
Method ModelNet10 (10) Omniobject3D (83)
Linear MLP-3 Linear MLP-3
MGM 77.3 83.1 50.3 57.5
+DINO 74.5 80.2 40.5 42.4
+iBOT 65.2 73.8 38.4 40.8
+LA 68.1 74.8 26.5 31.2
MGM+LA 83.1 84.6 47.3 53.2
Table H: Cross Domain Linear Probe Experiments. We report the mAcc results in the ablation of GaussianSSL on object-level classification tasks using the linear probe.
To evaluate global scene understanding, we design a cross-domain classification task (Tab. H) using object-level Gaussian splats from [28, 29]. We freeze the encoder and embedding layers and train only (1) a linear probe (final layer) to map the features to logits. (2) a 3-layer MLP (512 → 256 → classes) for nonlinear evaluation. We find that progressively adding DINO, iBOT, and language alignment (LA) losses degrades the classification accuracy. This misaligned trend with scene-level benchmarks stems from domain gaps, where the model must map distinct objects (e.g., fridge, oven) to similar scene-level semantics (e.g., ”kitchen”). Using masked pretraining and language alignment loss achieves the best performance on ModelNet10 (furniture), whereas MGM alone excels on OmniObject3D (common objects). For ModelNet10, MGM achieves the best 80% accuracy, whereas for the challenging object dataset (e.g., OmniObject3D), it peaks at  50%.
C Datasets Curation and Statistics
C.1 Analysis
Figure F: Example Incomplete Scenes in ARKitScenes.
Figure G: Blurry Scenes from 3RScan.
C.2 Visualization
Hypersim GS. Ground truth (left) and 3DGS rendering results (right).
D Dataset License
@@
SceneSplat-7K is constructed from several established indoor datasets, each attached with a specific license: ARKitScenes [1] (Apple Open Source License), Replica [45] (Replica Research License), ScanNet [5] (ScanNet Terms of Use), ScanNet++ [57] (ScanNet++ Terms of Use), Hypersim [42] (CC BY-SA 3.0), 3RScan [48] (3RScan Terms of Use), and Matterport3D [2] (Matterport Academic License Agreement). We have carefully structured our distribution approach to respect all original licenses while making our dataset accessible to the research community. For sources that permit redistribution, we plan to release our data on the Hugging Face under their respective terms of use. For datasets that require special permission, we co-host the data only after receiving approval from the original teams or let the original dataset team host the 3DGS data.
E Limitations